OpenStudy (anonymous):

A ball is thrown upward from a height of 50 feet with an initial velocity of 32 feet per second. The equation h=-16t^2+32t+50 gives the height h after t seconds. How long does it take for the ball to reach it's maximum height?

OpenStudy (anonymous):

y = -16x^2 + 32t + 50 Ball reaches max when t = -b/2a = -32/-32 = 1 sec